Grade 12Mechanics

Q.What is the moment of inertia of a cube about its main diagonal and how to find it?

sunindra  kanghujam
14 Years agoGrade 12

3 Answers

Aman  Bansal
14 Years ago

Dear Sunindra

FIrst lets find moment Jz of a³ cube around its z-axis:

dJz = r² dm = (x² + y²) ρ dx dy dz
dJz = (x² + y²) aρ dx dy
Jz = integral (x² + y²) aρ dx dy
Jz = 2 integral x² aρ dx dy
Jz = 2a²ρ integral x² dx
Jz = 2/3 a²ρ [x³ from -a/2 to a/2]
Jz = 1/6 ma²

Since Jx=Jx=Jz, and x,y,z are obviously main axes of
tensor Jij of inertia, the elliopsoid of inertia of a cube
is perfectly degerate spherical.

Answer:
Moment of inertia of a cube around any axis
passing throug it center of mass is ma²/6

Best Of Luck...!!!!

Cracking IIT just got more exciting,It’s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and winexciting gifts by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple click here to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert

Gravemind
8 Years ago
Divide cube into 8 equal cubes. Let moment of inertia of each be I. Then moment of inertia of bigger cube is 32I since mass becomes 8 times and side length becomes 2 times. Applying parallel axes theorem, we get the equation 32I = 8I + 6(2/3 ML^2) solving which, we get I = ML^2/6 , which is the required value.
ankit singh
6 Years ago
ivide cube into 8 equal cubes. Let moment of inertia of each be I. Then moment of inertia of bigger cube is 32I since mass becomes 8 times and side length becomes 2 times. Applying parallel axes theorem, we get the equation 32I = 8I + 6(2/3 ML^2) solving which, we get I = ML^2/6 , which is the required value.